OpenStudy (anonymous):

Find the inverse f(x)=x-7

OpenStudy (anonymous):

=7

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y=x-7 x=y-7 y=x+7

OpenStudy (anonymous):

x+7

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

F-1(x) = x+7